# Drell–Yan at the Electron-Ion Collider

### v1

@@ I Introduction
where the initial-state photon is supplied by the electron beam. A leading-order diagram contributing to this process is sketched in Fig. 1, where $X_{\gamma}$ and $X_{p}$ represent the hadronic remnants of the photon and proton, respectively. The power of measurements of di-jet photoproduction at the EIC to constrain the polarized and unpolarized PDFs of the photon has been demonstrated in Ref. Chu:2017mnm. The hadron-beam Drell–Yan process, $h+h\rightarrow\ell^{+}\ell^{-}+X$, is a primary tool in the phenomenology of both collinear and TMD parton distributions222“Drell–Yan” is often used colloquially to refer to the quark-antiquark annihilation subprocess, but other subprocesses such as QCD Compton scattering ($q+g\rightarrow q+\gamma^{*}$) also produce virtual photons which decay to a dilepton pair. The subprocess which produced the virtual photon cannot be distinguished experimentally, so the various subprocesses need to be calculated and included at the level of phenomenological analyses. For the purposes of this paper we will refer to all QCD virtual-photon-producing processes as Drell–Yan.. For reviews, see Refs. Reimer:2007Review; Peng:2014hta. In addition to $p+p$ collisions, experimental measurements of $\pi/K+p\rightarrow\ell^{+}\ell^{-}+X$ on polarized and unpolarized proton targets have provided novel constraints on the partonic structure of both beam and target COMPASS2017DYTSA; COMPASS:2024PRL; Peng:2017KaonJpsi; Bourrely:2024KaonPDF; E615:1989PRD; Barry:2018PRL; Novikov:2020xFitter; NA10:1988ZPC.
@@ II.2 Kinematic Reconstruction
One unique aspect of this program at the EIC is the ability to identify hadron species in jets. For resolved-photon events, which nominally arise from a $q\bar{q}$ configuration of the photon, a parton of a given flavor participates in the hard scattering, while the partner of that parton forms the photon remnant. Identifying the hadron species inside the remnant jet in the photon hemisphere can provide flavor information on the struck parton in the photon PDF. The capabilities of jet flavor identification in photoproduction have already been explored in considerable detail in Ref. Chu:2017mnm, to which we refer the reader for a more comprehensive discussion. This is also conceptually similar to the HERA measurements of heavy flavor di-jets in resolved photoproduction H1:2012oud; ZEUS:2011aa. In the $\gamma p$ center-of-mass frame the remnant is emitted predominantly into the photon hemisphere, i.e. $y^{\ast}\!\ll\!0$, where $y^{\ast}$ refers to the rapidity in the $\gamma p$ center-of-mass frame as opposed to the lab frame. The amount of energy available for the photon remnant is $E_{\gamma}(1-x_{\gamma})$. Thus, if $E_{\gamma}$ is known from the tagged scattered electron, the remnant jet energy can therefore be used as an alternative method to reconstruct $x_{\gamma}$. One limitation is that the momentum available for the remnant can become small when $y$ is small and $x_{\gamma}$ is large. Most likely a special jet reconstruction algorithm which takes into account the kinematic information of the event (à la Centauro Arratia:2020ssx) will be required to reconstruct jets this case. After applying detector effects, the resolution on the momentum of the hadronic final state in the remnant hemisphere is around 300 MeV and is roughly constant with energy. Therefore at small $y$ and large $x_{\gamma}$ the resolution on $x_{\gamma}$ from the remnant method degrades significantly.
@@
Using the expected tracking resolutions of the ePIC detector, the resolutions on the lepton pair mass, transverse momentum, and $\gamma p$ center-of-mass frame rapidity are shown in Fig. 6. The final product of the analysis would be an inclusive dilepton spectrum similar to the one shown in Fig. 8, or in Refs. PHENIX:2009gyd; PHENIX:2018dwt, from which the Drell–Yan contribution can be extracted after a careful background subtraction. The response matrix for $x_{\gamma}^{\mathrm{LO}}=\frac{M_{\ell\ell}}{\sqrt{s_{\gamma p}}}e^{-y_{\ell\ell}}$ in tagged photoproduction is shown in Fig. 4. The resolution on $x_{\gamma}^{\mathrm{LO}}$ is somewhat better in Drell-Yan than in the previous results for dijet production (See Fig. 4 of Ref. Chu:2017mnm). This is a result of the facts that the Low-$Q^{2}$ tagger is now a tracking system with better resolution than the calorimeter that was assumed in Ref. Chu:2017mnm, and the kinematic reconstruction in Drell-Yan relies only on single charged particles, which are generally less susceptible to large resolution fluctuations than jets.
@@
Triggering is a non-trivial aspect of this measurement because the signal is embedded in an enormous rate of inclusive photoproduction events. At HERA, photoproduction triggers were therefore typically heavily prescaled. For the dielectron channel, $\gamma p\to e^{+}e^{-}+X$, the signature of an $e^{+}e^{-}$ pair in the central detector should be relatively easy to trigger on, even for a more traditional trigger system. Dimuon selections, however, were practical at HERA only because dedicated muon systems provided a relatively clean low-level signal to trigger on, but such muon systems are currently not foreseen for ePIC. At the EIC, the intent is to use streaming readout with no hardware trigger, so the keep/reject decision can be deferred until after more advanced online reconstruction and feature building (e.g. cluster-track matching) have been performed AbdulKhalek:2021gbh; Ameli:2022gvw; Brei:2025isx. Requiring two tracks pointing to MIP-like energy deposits throughout the calorimeters will significantly reduce the background rate. If a hadron rejection factor of 50 can be obtained online for each muon, it would reduce the rate of events being read out by a factor of 2500 compared to a simple trigger on two tracks. This means that streaming readout should allow the dimuon channel to be retained with higher efficiency even in inclusive photoproduction, where a traditional hardware trigger would likely be forced into aggressive prescales or thresholds. For the remainder of this paper, we assume that the topology of $e^{\prime}$ in the low-$Q^{2}$ tagger and $\ell^{+}\ell^{-}$ in the central detector can be accepted with good efficiency. In reality, the feasibility of accepting the dimuon channel will ultimately depend on the final implementation of the data acquisition system.
@@ II.3 Backgrounds
The largest physics background for Drell–Yan at relatively low masses ($M_{\ell\ell}<4$ GeV) arises from decays of hadrons to leptons PHENIX:2009gyd; PHENIX:2018dwt. For this reason, experiments targeting the Drell–Yan process historically have considered masses greater than the $J/\psi$ mass. However, the excellent tracking resolution of the EIC means that peaking sources of dileptons such as the $J/\psi$ should be very narrow and can be subtracted from the continuum. Measurements of low-mass dielectrons and dimuons have been performed in $p+p$ collisions for example by PHENIX in Refs. PHENIX:2009gyd; PHENIX:2018dwt, where it was observed that in the mass range between the $\phi$ and $J/\psi$ mesons the $M_{\ell\ell}$ spectrum was dominated by semi-leptonic decays of charmed hadrons arising from correlated $c\bar{c}$ pairs. This background, in principle, can be suppressed at the EIC with precise vertexing. The removal of displaced vertices from semi-leptonic weak decays of $D$-mesons and other charm hadrons can help reduce the background sources of dilepton pairs in the continuum region of 1-3 GeV.
@@
There are also various physics processes which produce the same continuum dilepton final state. One of these is the Bethe–Heitler reaction shown in the top panel of Fig. 5, whereby the incoming photon interacts with a photon originating from the proton and produces a lepton pair Chwastowski:2022fzk. One can also imagine a dilepton background contribution from the process where the electron scatters on a positively charged lepton from the lepton PDF of the proton, as discussed in Ref. DaRold:2024ram. The primary background to Drell-Yan arises from the case where the photon splits into a dilepton pair, and one of the leptons emits a virtual photon which scatters inelastically on a parton in the proton and produces two jets, a target remnant and a current jet. This process is shown in the top left of Fig. IV. The elastic or quasi-elastic case where the proton remains intact or dissociates into a low-mass state can be easily distinguished and subtracted via the pattern of produced final-state hadrons. Although the Drell–Yan and inelastic Bethe–Heitler amplitudes formally interfere (a potentially useful feature that will be discussed in Sec. IV), the two contributions populate rather different regions of phase space, so that one can select kinematics where the cross section is dominated by Drell–Yan and the Bethe–Heitler background is small Busenitz:1981hr. Bethe–Heitler events reside almost entirely at $x_{\gamma}\simeq 1$ and very small $|t|$, with the lepton pair boosted into the photon hemisphere, whereas resolved-photon Drell–Yan populates the full region $x_{\gamma}<1$ and has a much flatter rapidity dependence Busenitz:1981hr. By requiring $x_{\gamma}<1$ and moderate or negative dilepton rapidity, the Bethe–Heitler contribution can be driven below the percent-level relative to Drell–Yan, as observed in Ref. Busenitz:1981hr. The angular distributions of the leptons in the in the dilepton rest frame can further discriminate Drell–Yan from purely QED production Bussey:1996vq, as the Bethe–Heitler contribution peaks very strongly at certain angles while Drell–Yan does not. The remaining contribution of Bethe–Heitler can be calculated and subtracted as a background.
@@
In Drell–Yan measurements, often the dimuon final state is the channel of choice for experimental measurements. At the EIC, both the dielectron and dimuon channels are available, but each has their own unique challenges. For dimuons, the challenge is cleanly identifying muons amongst a large pion background. The yield of pions in minimum bias $ep$ events is approximately a factor of 500 higher than the yield of muons. The desired pion rejection factor to reconstruct dimuon events with 90% efficiency and 10% contamination is around 5000. Unlike most fixed-target Drell–Yan or $pp$ collider experiments, ePIC does not have dedicated muon identification detectors. A cut on the energy deposited in the calorimeters can provide a pion rejection factor of about 100, which is sufficient for reconstructing vector mesons whose yields can be extracted from invariant-mass peaks above background, but for continuum processes such as Drell–Yan a higher muon purity is needed. Depending on the muon momentum, the time-of-flight and Cherenkov PID systems will provide additional muon–pion separation; however, the combined performance is unlikely to reach the level required for a clean dimuon measurement [^7] Comprehensive studies of the muon identification capabilities of ePIC are currently ongoing.. Such a measurement may therefore be better suited to a second EIC detector optimized for high-purity muon detection Jacobs:2025rqa.
@@
For the dielectron channel, a DIS scattered electron, particularly if it radiates a photon in the initial-state, can be misidentified as a Drell–Yan pair electron and erroneously paired with a positron from a hadronic decay. Since the DIS cross section is much larger than the Drell–Yan cross section and also will produce a continuum shape, this background is an added complication for the dielectron measurement. With the inclusion of tagging the scattered electron in the Low-$Q^{2}$ tagger, this background should be reduced. In contrast to the dimuon channel, the purity of electrons in ePIC will be very high, particularly in the region of $\eta<1.4$ corresponding to the barrel and backward regions where the combined hadron rejection power of particle identification and calorimetry is expected to be on the order of 104. Around two-thirds of the Drell–Yan leptons have $\eta<1.4$, and this region generally corresponds to larger values of $x_{\gamma}$. Another downside of the dielectron channel is the increased magnitude of radiative corrections.
@@
One unique opportunity at the EIC for relatively large $M_{\ell\ell}$ is to exploit the $\gamma^{*}\rightarrow\tau^{+}\tau^{-}$ channel. Reconstruction of $\tau$ leptons has been discussed at the EIC in the context of charged-lepton flavor violation and measuring the $\tau$ anomalous electromagnetic moment Zhang:2023EICetau; Deng:2025hio; Chwastowski:2022fzk. While the number of usable di-tau events will be significantly lower than dielectrons or dimuons due to the unfavorable branching fractions of the $\tau$, the di-tau channel has the benefit that the displaced vertex of the $\tau$ can dramatically reduce the background. The combination of three dilepton final-states, each of which have different sources of background, will provide a cross-check and improve the understanding of the Drell–Yan component of the inclusive dilepton spectrum.
@@ III.1 Cross Sections
One of the fortuitous observations is that for the 10x275 GeV beam energy configuration, the majority of the Drell–Yan dilepton pairs enter the central detector acceptance, as can be seen from Fig. 3. The reason is that the photon PDF prefers larger values of $x$ than the proton PDF, but the energy of the proton beam is much higher than the photon. The result is that the quark-antiquark collisions are not dramatically boosted in either the proton-going or electron-going directions in the lab-frame. The center panel of Fig. 3 shows that the detector acceptance in the forward direction does not significantly limit the reach to large $x_{\gamma}$, but the reach to low-$x_{\gamma}$ is somewhat limited by the acceptance in the backward direction.
Figure 6: Expected reconstructed event yields for tagged photoproduction at the EIC for the SAS1M and SAS2M photon PDF sets. The yields are shown for a single dilepton channel, i.e. they would be doubled if including both dielectron and dimuon channels. Detector acceptances and resolutions are imposed on the scattered electron and dilepton pair. The lower panels provide the expected detector resolutions on the variables given the detector parameters in the Appendix A.
@@
The dielectron invariant mass distribution from Drell–Yan, $J/\psi\mathrm{~and~}\Upsilon$ production is shown in Fig. 8. The dielectron channel is chosen since electrons will likely have a higher purity than muons, at the price of larger radiative corrections. The inelastic $J/\psi$ cross section was measured by H1 H1:2010udv and ZEUS ZEUS:1997wrc; ZEUS:2002src to be approximately 2 nanobarns for the range of $W$ relevant to tagged photoproduction at the EIC. The $\psi^{\prime}$ cross section was measured by ZEUS to be a factor of 5 or so smaller than the $J/\psi$ cross section. The yield of $\Upsilon$ decaying to the dielectron channel is roughly approximated from the elastic HERA data of Refs. ZEUS:2009Upsilon, assuming the elastic and inelastic cross sections are similar, as is the case for $J/\psi$. The cross sections for the $\Upsilon(2\mathrm{S})$ and $\Upsilon(3\mathrm{S})$ are assumed to be 10% and 5% of the $\Upsilon(1\mathrm{S})$, respectively. The shape of the $M_{\ell\ell}$ spectrum in Fig. 8 includes the expected resolutions of ePIC. The binning is chosen such that at all points the bin width is at least two times the resolution on $M_{\ell\ell}$. Combinatorial and misidentification backgrounds are excluded, since the degree to which they can be subtracted will depend on the final detector performance.
@@
Figure 9: Expected Bjorken $x$ distributions for the proton and photon in tagged Drell–Yan events at the EIC. The darker colored curves show the effect of cutting on $M_{\ell\ell}>M_{J/\psi}$.

### v2

@@ I Introduction
where the initial-state photon is supplied by the electron beam. A leading-order diagram contributing to this process is sketched in Fig. 1, where $X_{\gamma}$ and $X_{p}$ represent the hadronic remnants of the photon and proton, respectively. The power of measurements of di-jet photoproduction at the EIC to constrain the polarized and unpolarized PDFs of the photon has been demonstrated in Ref. Chu:2017mnm. The hadron-beam Drell–Yan process, $h+h\rightarrow\ell^{+}\ell^{-}+X$, is a primary tool in the phenomenology of both collinear and TMD parton distributions222“Drell–Yan” is often used colloquially to refer to the quark-antiquark annihilation subprocess, but other subprocesses such as QCD Compton scattering ($q+g\rightarrow q+\gamma^{*}$) also produce virtual photons which decay to a dilepton pair. The subprocess which produced the virtual photon cannot be distinguished experimentally, so the various subprocesses need to be calculated and included at the level of phenomenological analyses. For the purposes of this paper we will refer to all QCD virtual-photon-producing inelastic processes as Drell–Yan.. For reviews, see Refs. Reimer:2007Review; Peng:2014hta. In addition to $p+p$ collisions, experimental measurements of $\pi/K+p\rightarrow\ell^{+}\ell^{-}+X$ on polarized and unpolarized proton targets have provided novel constraints on the partonic structure of both beam and target COMPASS2017DYTSA; COMPASS:2024PRL; Peng:2017KaonJpsi; Bourrely:2024KaonPDF; E615:1989PRD; Barry:2018PRL; Novikov:2020xFitter; NA10:1988ZPC.
@@ II.2 Kinematic Reconstruction
One unique aspect of this program at the EIC is the ability to identify hadron species in jets. For resolved-photon events, which nominally arise from a $q\bar{q}$ configuration of the photon, a parton of a given flavor participates in the hard scattering, while the partner of that parton forms the photon remnant. Identifying the hadron species inside the remnant jet in the photon hemisphere can provide flavor information on the struck parton in the photon PDF. The capabilities of jet flavor identification in photoproduction have already been explored in considerable detail in Ref. Chu:2017mnm, to which we refer the reader for a more comprehensive discussion. This method of PDF flavor separation is also conceptually similar to the HERA measurements of heavy flavor di-jets in resolved photoproduction H1:2012oud; ZEUS:2011aa. In the $\gamma p$ center-of-mass frame, the remnant is emitted predominantly into the photon hemisphere, i.e., $y^{\ast}\!\ll\!0$, where $y^{\ast}$ refers to the rapidity in the $\gamma p$ center-of-mass frame as opposed to the lab frame. The amount of energy available for the photon remnant is $E_{\gamma}(1-x_{\gamma})$. Thus, if $E_{\gamma}$ is known from the tagged scattered electron, the remnant jet energy can therefore be used as an alternative method to reconstruct $x_{\gamma}$. One limitation is that the momentum available for the remnant can become small when $y$ is small and $x_{\gamma}$ is large. Most likely a special jet reconstruction algorithm which takes into account the kinematic information of the event (à la Centauro Arratia:2020ssx) will be required to reconstruct jets in this case. After applying detector effects, the resolution on the momentum of the hadronic final state in the remnant hemisphere is around 300 MeV and is roughly constant with energy. Therefore, at small $y$ and large $x_{\gamma}$ the resolution on $x_{\gamma}$ from the remnant method degrades significantly.
@@
Using the expected tracking resolutions of the ePIC detector, the resolutions on the lepton pair mass, transverse momentum, and $\gamma p$ center-of-mass frame rapidity are shown in Fig. 6. The final product of the analysis would be an inclusive dilepton spectrum similar to the one shown in Fig. 8, or in Refs. PHENIX:2009gyd; PHENIX:2018dwt, from which the Drell–Yan contribution can be extracted after a careful background subtraction. The response matrix for $x_{\gamma}^{\mathrm{LO}}=\frac{M_{\ell\ell}}{\sqrt{s_{\gamma p}}}e^{-y_{\ell\ell}}$ in tagged photoproduction is shown in Fig. 4. The resolution on $x_{\gamma}^{\mathrm{LO}}$ is somewhat better in Drell-Yan than in the previous results for dijet production (See Fig. 4 of Ref. Chu:2017mnm). This improvement arises from the fact that the Low-$Q^{2}$ tagger is now a tracking system with better resolution than the calorimeter that was assumed in Ref. Chu:2017mnm and because the kinematic reconstruction in Drell-Yan relies only on single charged particles, which are generally less susceptible to large resolution fluctuations than jets.
@@
Triggering is a non-trivial aspect of this measurement because the signal is embedded in an enormous rate of inclusive photoproduction events. At HERA, photoproduction triggers were therefore typically heavily prescaled. For the dielectron channel, $\gamma p\to e^{+}e^{-}+X$, the signature of an $e^{+}e^{-}$ pair in the central detector should be relatively easy to trigger on, even for a more traditional trigger system. Dimuon selections, however, were practical at HERA only because dedicated muon systems provided a relatively clean low-level signal to trigger on, but such muon systems are currently not foreseen for ePIC. At the EIC, the intent is to use streaming readout with no hardware trigger, so the keep/reject decision can be deferred until after more advanced online reconstruction and feature building (e.g., cluster-track matching) have been performed AbdulKhalek:2021gbh; Ameli:2022gvw; Brei:2025isx. Requiring two tracks pointing to MIP-like energy deposits throughout the calorimeters will significantly reduce the background rate. If a hadron rejection factor of 50 can be obtained online for each muon, it would reduce the rate of events being read out by a factor of 2500 compared to a simple trigger on two tracks. This means that streaming readout should allow the dimuon channel to be retained with higher efficiency even in inclusive photoproduction, where a traditional hardware trigger would likely be forced into aggressive prescales or thresholds. For the remainder of this paper, we assume that the topology of $e^{\prime}$ in the low-$Q^{2}$ tagger and $\ell^{+}\ell^{-}$ in the central detector can be accepted with good efficiency. In reality, the feasibility of accepting the dimuon channel will ultimately depend on the final implementation of the data acquisition system.
@@ II.3 Backgrounds
The largest physics background for Drell–Yan at relatively low masses ($M_{\ell\ell}<4$ GeV) arises from decays of hadrons to leptons PHENIX:2009gyd; PHENIX:2018dwt. For this reason, experiments targeting the Drell–Yan process historically have considered masses greater than the $J/\psi$ mass. However, the excellent tracking resolution of the EIC means that peaking sources of dileptons such as the $J/\psi$ should be very narrow and can be subtracted from the continuum. Measurements of low-mass dielectrons and dimuons have been performed in $p+p$ collisions for example by PHENIX in Refs. PHENIX:2009gyd; PHENIX:2018dwt, where it was observed that in the mass range between the $\phi$ and $J/\psi$ mesons, the $M_{\ell\ell}$ spectrum was dominated by semi-leptonic decays of charmed hadrons arising from correlated $c\bar{c}$ pairs. This background, in principle, can be suppressed at the EIC with precise vertexing. The removal of displaced vertices from semi-leptonic weak decays of $D$-mesons and other charm hadrons can help reduce the background sources of dilepton pairs in the continuum region of 1-3 GeV.
@@
There are also various physics processes which produce the same continuum dilepton final state. One of these is the Bethe–Heitler reaction shown in the top panel of Fig. 5, whereby the incoming photon interacts with a photon originating from the proton and produces a lepton pair Chwastowski:2022fzk. One can also imagine a dilepton background contribution from the process where the electron scatters on a positively charged lepton from the lepton PDF of the proton, as discussed in Ref. DaRold:2024ram. The primary background to Drell-Yan arises from the case where the photon splits into a dilepton pair, and one of the leptons emits a virtual photon which scatters inelastically on a parton in the proton and produces two jets: a target remnant and a current jet. This process is shown in the top left of Fig. IV. The elastic or quasi-elastic case where the proton remains intact or dissociates into a low-mass state can be easily distinguished and subtracted via the pattern of produced final-state hadrons. Although the Drell–Yan and inelastic Bethe–Heitler amplitudes formally interfere (a potentially useful feature that will be discussed in Sec. IV), the two contributions populate rather different regions of phase space, so that one can select kinematics where the cross section is dominated by Drell–Yan and the Bethe–Heitler background is small Busenitz:1981hr. Bethe–Heitler events reside almost entirely at $x_{\gamma}\simeq 1$ and very small $|t|$, with the lepton pair boosted into the photon hemisphere, whereas resolved-photon Drell–Yan populates the full region $x_{\gamma}<1$ and has a much flatter rapidity dependence Busenitz:1981hr. By requiring $x_{\gamma}<1$ and moderate or negative dilepton rapidity, the Bethe–Heitler contribution can be driven below the percent-level relative to Drell–Yan, as observed in Ref. Busenitz:1981hr. The angular distributions of the leptons in the dilepton rest frame can further discriminate Drell–Yan from purely QED production Bussey:1996vq, as the Bethe–Heitler contribution peaks very strongly at certain angles while Drell–Yan does not. The remaining contribution of Bethe–Heitler can be calculated and subtracted as a background.
@@
In Drell–Yan measurements, often the dimuon final state is the channel of choice for experimental measurements. At the EIC, both the dielectron and dimuon channels are available, but each has their own unique challenges. For dimuons, the challenge is cleanly identifying muons amongst a large pion background. The yield of pions in minimum bias $ep$ events is approximately a factor of 500 higher than the yield of muons. The desired pion rejection factor to reconstruct dimuon events with 90% efficiency and 10% contamination is around 5000. Unlike most fixed-target Drell–Yan or $pp$ collider experiments, ePIC does not have dedicated muon identification detectors. A cut on the energy deposited in the calorimeters can provide a pion rejection factor of about 100, which is sufficient for reconstructing vector mesons whose yields can be extracted from invariant-mass peaks above background, but for continuum processes such as Drell–Yan, a higher muon purity is needed. Depending on the muon momentum, the time-of-flight and Cherenkov PID systems will provide additional muon–pion separation; however, the combined performance is unlikely to reach the level required for a clean dimuon measurement [^7] Comprehensive studies of the muon identification capabilities of ePIC are currently ongoing.. Such a measurement may therefore be better suited to a second EIC detector optimized for high-purity muon detection Jacobs:2025rqa.
@@
For the dielectron channel, a DIS scattered electron, particularly if it radiates a photon in the initial-state, can be misidentified as a Drell–Yan pair electron and erroneously paired with a positron from a hadronic decay. Since the DIS cross section is much larger than the Drell–Yan cross section and also will produce a continuum shape, this background is an added complication for the dielectron measurement. With the inclusion of tagging the scattered electron in the Low-$Q^{2}$ tagger, however, this background should be reduced. In contrast to the dimuon channel, the purity of electrons in ePIC will be very high, particularly in the region of $\eta<1.4$ corresponding to the barrel and backward regions where the combined hadron rejection power of particle identification and calorimetry is expected to be on the order of 104. Around two-thirds of the Drell–Yan leptons have $\eta<1.4$, and this region generally corresponds to larger values of $x_{\gamma}$. Another downside of the dielectron channel is the increased magnitude of radiative corrections.
@@
One unique opportunity at the EIC for relatively large $M_{\ell\ell}$ is to exploit the $\gamma^{*}\rightarrow\tau^{+}\tau^{-}$ channel. Reconstruction of $\tau$ leptons has been discussed at the EIC in the context of charged-lepton flavor violation and measuring the $\tau$ anomalous electromagnetic moment Zhang:2023EICetau; Deng:2025hio; Chwastowski:2022fzk. While the number of usable di-tau events will be significantly lower than dielectrons or dimuons due to unfavorable branching fractions, the di-tau channel provides a dramatically reduced background due to the displaced vertex of the $\tau$. The combination of three dilepton final-states, each of which have different sources of background, will provide a cross-check and improve the understanding of the Drell–Yan component of the inclusive dilepton spectrum.
@@ III.1 Cross Sections
One of the fortuitous observations is that for the 10x275 GeV beam energy configuration, the majority of the Drell–Yan dilepton pairs enter the central detector acceptance, as can be seen from Fig. 3. This occurs because the photon PDF prefers larger values of $x$ than the proton PDF, but the energy of the proton beam is much higher than the photon. The result is that the quark-antiquark collisions are not dramatically boosted in either the proton-going or electron-going directions in the lab-frame. The center panel of Fig. 3 shows that the detector acceptance in the forward direction does not significantly limit the reach to large $x_{\gamma}$, but the reach to low-$x_{\gamma}$ is somewhat limited by the acceptance in the backward direction.
Figure 6: Expected reconstructed event yields for tagged photoproduction at the EIC for the SAS1M and SAS2M photon PDF sets. The yields are shown for a single dilepton channel, i.e., they would be doubled if including both dielectron and dimuon channels. Detector acceptances and resolutions are imposed on the scattered electron and dilepton pair. The lower panels provide the expected detector resolutions on the variables given the detector parameters in Appendix A.
@@
The dielectron invariant mass distribution from Drell–Yan, $J/\psi\mathrm{,~and~}\Upsilon$ production is shown in Fig. 8. The dielectron channel is chosen since electrons will likely have a higher purity than muons, at the price of larger radiative corrections. The inelastic $J/\psi$ cross section was measured by H1 H1:2010udv and ZEUS ZEUS:1997wrc; ZEUS:2002src to be approximately 2 nanobarns for the range of $W$ relevant to tagged photoproduction at the EIC. The $\psi^{\prime}$ cross section was measured by ZEUS to be a factor of 5 or so smaller than the $J/\psi$ cross section. The yield of $\Upsilon$ decaying to the dielectron channel is roughly approximated from the elastic HERA data of Refs. ZEUS:2009Upsilon, assuming the elastic and inelastic cross sections are similar, as is the case for $J/\psi$. The cross sections for the $\Upsilon(2\mathrm{S})$ and $\Upsilon(3\mathrm{S})$ are assumed to be 10% and 5% of the $\Upsilon(1\mathrm{S})$, respectively. The shape of the $M_{\ell\ell}$ spectrum in Fig. 8 includes the expected resolutions of ePIC. The binning is chosen such that at all points the bin width is at least two times the resolution on $M_{\ell\ell}$. Combinatorial and misidentification backgrounds are excluded, since the degree to which they can be subtracted will depend on the final detector performance.
@@
Figure 9: Expected Bjorken $x$ distributions for the proton and photon in tagged Drell–Yan events at the EIC. The darker–colored curves show the effect of cutting on $M_{\ell\ell}>M_{J/\psi}$.